# ReAttention: Training-Free Infinite Context with Finite Attention Scope

### v2

@@ 3.4 Efficiency Analysis
We analyze our kernel’s efficiency in terms of GPU execution time and memory usage, showing that using Triton rather than PyTorch for the top-$k$ attention greatly enhances GPU performance while reducing GPU memory consumption. Additionally, we compare ReAttention with the standard implementation in the HuggingFace Transformers(Dai et al., 2019), measuring time to first token (TTFT) and memory cost in context lengths 32k to 256k. All experiments were conducted on a system with a 48-core CPU, 256GB RAM, and an A800-80GB GPU.
@@
Memory Usage in Prefilling Stage
Memory overhead limits the prefilling for longer sequences. Figure 7 shows that our method surpasses the standard implementation in HuggingFace Transformers as context length grows, continuing to work when the standard implementation runs out of memory. We also recorded first-token latency, which remains comparable to the standard implementation.
@@ 4.1 Analysis on Hyper-parameter
We find that as the chunk size decreases from 2048 to 512, the performance gradually improves. The prefilling stage becomes more precise in extracting critical information during each top-$k$ attention step. However, since prefilling takes more time as chunk size decreases, we select 512 as our default value without further reduction. Next, we examined the impact of different span sizes. Among these hyper-parameters, span size has the most significant effect on downstream performance. When span size is relatively small, fragmented segments weaken coherence and may mislead the LLM’s predictions, resulting in poorer performance compared to StreamingLLM. When the span size reaches 32, the performance on LongBench peaks. Beyond this point, due to the constraint of pre-trained position embedding, the number of extracted segments decreases, hindering the effective capture of critical contextual information. We also compare the influence of the top-$k$ number and find that top-4 yields the best results. A smaller top-$k$ number makes it difficult for LLM to identify critical information, while a larger top-$k$ number may introduce irrelevant content that interferes with judgment. Finally, we analyze the effect of local size. Smaller local size compromises LLM’s ability to maintain semantic coherence, and excessive selection beyond the local context may skew the KV cache during the prefilling stage. Therefore, we set the local size to half of the pre-training context length for LLaMA3-8B-8K and apply this hyper-parameter configuration to other LLMs.
@@ 4.2 Postion-Agnostic Cache Selection
Figure 8: Visualization of the attention distributions for InternLM2-7B-200K(Cai et al., 2024) evaluated on Needle-In-A-Haystack(Contributors, 2023a; b) in 32K context length. In each sub-figure, each heatmap represents the attention distribution of a single layer, with the y-axis corresponding to the 32 attention heads and the x-axis corresponding to the 32K context length. The first 32 tokens, the instruction part, and the last 6 tokens, the generated part, are not excluded. The values represent the cumulative self-attention distributions for each token generation. The brighter the color, the higher the cumulative attention. A pooling operation with a kernel size of 100 is applied sequentially to show the pattern more clearly. The red box represents the position of the ”needle” in the context.
In ReAttention, the full-context cache selection is based on the dot product between the query and key vectors without position embedding. This may create a gap between the selection of the KV cache and the self-attention distribution. Therefore, here raises the question of whether using semantic vectors without position information can effectively locate the critical information in the context. To analyze this, we select a correct case and a wrong case from the results of Needle-In-A-Haystack evaluation(Contributors, 2023a; b) on InternLM2-7B-200K(Cai et al., 2024) with full attention in 32K context length. For each case, we calculate both the attention distribution with position embedding, i.e., the real self-attention distribution, and the attention distribution without position embedding.
@@ 4.3 Observations on Synthetic Tasks
The above analysis has demonstrated ReAttention’s good performance in various long-context scenarios, suggesting that it can extrapolate to infinite context lengths. However, these evaluations primarily focus on common natural contexts and lack currently widely-discussed benchmarks(Kuratov et al., 2024; Li et al., 2024), such as RULER(Hsieh et al., 2024). Unfortunately, ReAttention performs poorly on the RULER benchmark. In fact, the extrapolation methods not based on full attention, including ReAttention and InfLLM, fail to pass the RULER benchmark. For instance, using the LLaMA3-8B-8K(Meta, 2024a), we compare Dynamic NTK extrapolation with InfLLM(Xiao et al., 2024) and ReAttention at 8K and 16K context lengths. The results are shown in Table 3. Only Dynamic NTK achieves effective extrapolation, while both ReAttention and InfLLM exhibit more pronounced performance degradation as the context length increases. Specifically, We focus on reporting the results of two subtasks from RULER: NIAH-Single3 and NIAH-MultiKey3(Hsieh et al., 2024). Both tasks involve extracting key information from a context that contains misleading items. The difference is that the context for Single3 consists of natural text, where the key information is a string of random alphanumeric characters, whereas MultiKey3 features a stack of key-value pairs composed of those strings. We find that while all three methods can succeed in the Single3 test, only full-attention-based Dynamic NTK can effectively handle MultiKey3.
@@
Nevertheless, the synthetic chaotic long texts are quite rare in practical scenarios(Lv et al., 2024). In the common natural long text benchmarks we have tested, ReAttention consistently demonstrates strong performance. Therefore, ReAttention still remains competitive for application in extending the context length of LLMs to infinity. We provide an in-depth discussion of the characteristics reflected in the QK sequence extension patterns between synthetic and natural texts in Appendix B.
@@ 5 Related Work
In contrast, another line of research has tried to extend the context length of models through sparse attention. Considering the self-attention distribution tends to focus on the global and local context, StreamingLLM and LM-Infinite proposed the $\Lambda$-shaped attention window to achieve almost unlimited input length(Xiao et al., 2023; Han et al., 2023). However, since the input is accompanied by discarding the previous context, it still fails to extend the context length(Dong et al., 2024). Based on StreamingLLM, InfLLM(Xiao et al., 2024) and LongHeads(Lu et al., 2024) attempt to extend context through chunkwise retrieval from the middle cache, but the semantic fragmentation and block representation issues also affect downstream task performance(Luo et al., 2024). Recently, MInference(Jiang et al., 2024) and RetrievalAttentionLiu et al. (2024) utilize a dynamic cache selection to achieve significant speedup, but not attempt context extrapolation. In contrast, we propose ReAttention, which regards the cache selection as the preceding attention before the ordinary self-attention and achieves an infinite context with a finite attention scope, making LLMs free from the challenge of length extrapolation. We also minimize the top-$k$ attention kernel-wise like FlashAttention in self-attention(Dao et al., 2022; Dao, 2023).
@@ 6 Conclusion
Appendix A Detailed Evaluation Results
@@
En. Zh. Code Avg. Close- Ended Open- Ended Avg.
LLaMA3-8B-8K* 33.91 18.98 69.80 35.25 46.58 2.66 17.30
+ DynamicNTK 34.27 19.95 69.89 35.48 46.11 2.79 17.23
+ StreamingLLM 29.48 13.22 70.27 32.49 36.52 1.36 13.08
+ ReAttention 33.42 18.74 70.38 35.03 44.94 2.43 16.60
LLaMA3.1-8B-128K 36.38 28.64 70.63 38.79 45.30 0.50 15.44
+ StreamingLLM 33.41 24.24 70.50 37.06 42.36 0.38 14.37
+ ReAttention 36.14 28.56 70.48 38.63 45.25 0.50 15.42
LLaMA3.1-70B-128K 46.00 32.92 74.89 46.02 60.78 0.15 20.36
+ StreamingLLM 41.07 29.60 75.06 43.20 55.32 0.16 18.54
+ ReAttention 46.09 32.99 75.26 46.15 61.11 0.18 20.49
LLaMA3.2-3B-128K 48.03 21.84 55.73 40.76 55.75 20.63 32.33
+ StreamingLLM 38.86 15.86 54.16 35.08 43.27 18.66 26.87
+ ReAttention 46.16 19.45 55.17 39.43 53.09 20.12 31.11
Mistral-v0.3-7B-32K 35.74 18.59 65.13 35.61 41.97 4.33 16.88
+ StreamingLLM 33.05 16.23 64.85 34.01 36.49 4.45 15.13
+ ReAttention 35.45 18.14 65.32 35.48 41.95 4.09 16.71
InternLM2.5-7B-1M 55.49 41.46 52.62 47.27 62.96 24.99 37.64
+ StreamingLLM 49.77 38.98 52.97 44.03 57.97 23.30 34.85
+ ReAttention 55.60 41.58 52.72 47.35 63.47 24.36 37.40
Qwen2-7B-128K 43.47 31.23 67.65 41.34 53.43 15.41 28.08
+ StreamingLLM 39.83 28.55 67.39 39.16 50.07 14.30 26.22
+ ReAttention 42.97 30.92 67.56 41.01 53.18 15.35 27.96
Qwen2-72B-128K 61.94 41.85 71.15 54.71 68.17 25.44 39.68
+ StreamingLLM 54.03 38.12 70.91 50.08 60.43 22.38 35.06
+ ReAttention 61.53 42.00 71.39 54.53 67.31 25.34 39.33
Qwen2-1B-32K 34.65 23.65 57.98 33.21 41.28 15.59 24.16
+ StreamingLLM 33.12 21.56 58.32 32.40 37.46 15.31 22.69
+ ReAttention 34.94 24.12 58.58 33.51 41.00 16.10 24.40
@@
Table 4 includes all the data previously shown in Figure 2, along with some additional data, providing a detailed comparison of various LLMs including LLaMA(Meta, 2024a; Dubey et al., 2024; Meta, 2024b), Mistral(mistralai, 2024), InternLM(Cai et al., 2024; InternLM, 2024), and Qwen(Yang et al., 2024) series across LongBench(Bai et al., 2023) and L-Eval(An et al., 2023).
Appendix B Analysis of Cache Feature
B.1 The Method for Obtaining t-SNE Results

### v3

@@ 3.4 Efficiency Analysis
We analyze our kernel’s efficiency in terms of GPU execution time and memory usage, showing that using Triton rather than PyTorch for the top-$k$ attention greatly enhances GPU performance while reducing GPU memory consumption. Additionally, we compare ReAttention with the standard implementation in the HuggingFace Transformers (Dai et al., 2019), measuring time to first token (TTFT) and memory cost in context lengths 32k to 256k. The TTFT and throughput comparison results are detailed in Appendix C. All experiments were conducted on a system with a 48-core CPU, 256GB RAM, and an A800-80GB GPU.
Triton Operator Efficiency
@@
Memory Usage in Prefilling Stage
Memory overhead limits the prefilling for longer sequences. Figure 6 shows that our method surpasses the standard implementation in HuggingFace Transformers
@@
as context length grows, continuing to work when the standard implementation runs out of memory. We also record first-token latency, which remains comparable to the standard implementation in Huggingface Transformers.
@@ 4.1 Analysis on Hyper-parameter
We find that as the chunk size decreases from 2048 to 512, the performance gradually improves. The prefilling stage becomes more precise in extracting critical information during each top-$k$ attention step. However, since prefilling takes more time as chunk size decreases, we select 512 as our default value without further reduction. Next, we examine the impact of different span sizes. Among these hyper-parameters, span size has the most significant effect on downstream performance. When span size is relatively small, fragmented segments weaken coherence and may mislead the LLM’s predictions, resulting in poorer performance compared to StreamingLLM. When the span size reaches 32, the performance on LongBench peaks. Beyond this point, due to the constraint of pre-trained position embedding, the number of extracted segments decreases, hindering the effective capture of critical contextual information. We also compare the influence of the top-$k$ number and find that top-4 yields the best results. A smaller top-$k$ number makes it difficult for LLM to identify critical information, while a larger top-$k$ number may introduce irrelevant content that interferes with judgment. Finally, we analyze the effect of local size. Smaller local size compromises LLM’s ability to maintain semantic coherence, and excessive selection beyond the local context may skew the KV cache during the prefilling stage. Therefore, we set the local size to half of the pre-training context length for LLaMA3-8B-8K and apply this hyper-parameter configuration to other LLMs.
@@ 4.2 Postion-Agnostic Cache Selection
Figure 7: Visualization of the attention distributions for InternLM2-7B-200K (Cai et al., 2024b) evaluated on Needle-In-A-Haystack (Contributors, 2023a; b) in 32K context length. In each sub-figure, each heatmap represents the attention distribution of a single layer, with the y-axis corresponding to the 32 attention heads and the x-axis corresponding to the 32K context length. The first 32 tokens, the instruction part, and the last 6 tokens, the generated part, are not excluded. The values represent the cumulative self-attention distributions for each token generation. The brighter the color, the higher the cumulative attention. A pooling operation with a kernel size of 100 is applied sequentially to show the pattern more clearly. The red box represents the position of the ”needle” in the context.
In ReAttention, the full-context cache selection is based on the dot product between the query and key vectors without position embedding. This may create a gap between the selection of the KV cache and the self-attention distribution. Therefore, here raises the question of whether using semantic vectors without position information can effectively locate the critical information in the context. To analyze this, we select a correct case and a wrong case from the results of Needle-In-A-Haystack evaluation (Contributors, 2023a; b) on InternLM2-7B-200K (Cai et al., 2024b) with full attention in 32K context length. For each case, we calculate the attention distribution with position embedding, i.e., the real self-attention distribution, and the attention distribution without position embedding.
@@ 4.3 Observations on Synthetic Tasks
The above analysis has demonstrated ReAttention’s good performance in various long-context scenarios, suggesting that it can extrapolate to infinite context lengths. However, these evaluations primarily focus on common natural contexts and lack currently widely-discussed benchmarks (Kuratov et al., 2024; Li et al., 2024), such as RULER (Hsieh et al., 2024). Unfortunately, ReAttention performs poorly on the RULER benchmark. In fact, the extrapolation methods not based on full attention, including ReAttention and InfLLM, fail to pass the RULER benchmark. For instance, using the LLaMA3-8B-8K (Meta, 2024a), we compare Dynamic NTK extrapolation with InfLLM (Xiao et al., 2024a) and ReAttention at 8K and 16K context lengths. The results are shown in Table 4. Only Dynamic NTK achieves effective extrapolation, while both ReAttention and InfLLM exhibit more pronounced performance degradation as the context length increases. Specifically, We focus on reporting the results of two subtasks from RULER: NIAH-Single3 and NIAH-MultiKey3 (Hsieh et al., 2024). Both tasks involve extracting key information from a context that contains misleading items. The difference is that the context for Single3 consists of natural text, where the key information is a string of random alphanumeric characters, whereas MultiKey3 features a stack of key-value pairs composed of those strings. We find that while all three methods can succeed in the Single3 test, only full-attention-based Dynamic NTK can effectively handle MultiKey3.
@@
Nevertheless, the synthetic chaotic long texts are quite rare in practical scenarios (Lv et al., 2024). In the common natural long text benchmarks we have tested, ReAttention consistently demonstrates strong performance. Therefore, ReAttention still remains competitive for application in extending the context length of LLMs to infinity. We provide an in-depth discussion of the characteristics reflected in the QK sequence extension patterns between synthetic and natural texts in Appendix E.
@@ 5 Related Work
In contrast, another line of research has tried to extend the context length of models through sparse attention. Considering the self-attention distribution tends to focus on the global and local context, StreamingLLM and LM-Infinite proposed the $\Lambda$-shaped attention window to achieve almost unlimited input length (Xiao et al., 2023; Han et al., 2023). However, since the input is accompanied by discarding the previous context, it still fails to extend the context length (Dong et al., 2024). Based on StreamingLLM, InfLLM (Xiao et al., 2024a) and LongHeads (Lu et al., 2024) attempt to extend context through chunkwise retrieval from the middle cache, but the semantic fragmentation and block representation issues also affect downstream task performance (Luo et al., 2024). Recently, MInference (Jiang et al., 2024) and RetrievalAttentionLiu et al. (2024a) utilize a dynamic cache selection to achieve significant speedup, but not attempt context extrapolation. In contrast, we propose ReAttention, which regards the cache selection as the preceding attention before the ordinary self-attention and achieves an infinite context with a finite attention scope, making LLMs free from the challenge of length extrapolation. We also minimize the top-$k$ attention kernel-wise like FlashAttention in self-attention (Dao et al., 2022; Dao, 2023). Additionally, since ReAttention is conducted based on cache selection, we also expand our discussion on cache optimization, especially on token eviction methods like H2O (Zhang et al., 2024d) and SnapKV, in Appendix D
@@ 6 Conclusion
Acknowledgement
This work is supported by the National Key Research and Development Program of China (No. U24B20181). We also appreciate the constructive comments from reviewers in the rebuttal and add discussion on methodology, effectiveness, efficiency, and related work in Appendix A to Appendix D.
Appendix A Pseudocode of ReAttention
In this section, we present the pseudocode for ReAttention’s prefilling and decoding phases. Our approach uses position-agnostic cache selection at each generation step, as shown in Section 2.1.
Input: The input token indices input_ids, number of tokens $L$, the number of starting tokens $l_{\text{global}}$, the number of ending tokens $l_{\text{local}}$, size of prefilling chunk $l_{\text{chunk}}$.
1 Initialize: $s=0$, $e=l_{\text{global}}+l_{\text{local}}$, past_key_values = None
2
3while $s<L$ do
4       $\triangleright$ Process next chunk
5       past_key_values = None
6       input_chunk = input_ids[:, $s$:min($e$, $L$)]
7       hidden_states = embedding(input_chunk)
8       $\triangleright$ Forward through each transformer layer
9       for $i\leftarrow 0$ to $N_{\text{layer}}-1$ do
10             $\triangleright$ Project to Q, K, V
11             query = q_proj(hidden_states)
12             key = k_proj(hidden_states)
13             value = v_proj(hidden_states)
14             $\triangleright$ Update KV cache for this layer
15             if $\mathrm{past\_key\_values}$ then
16                   key = concat(past_key_values[$i$].key, key)
17                   value = concat(past_key_values[$i$].value, value)
18
19            past_key_values[$i$] = (key, value)
20             $\triangleright$ Position-agnostic cache selection
21             score = query @ key[$l_{\text{global}}$:-$l_{\text{local}}$]${}^{\texttt{T}}$
22             idx_recall = topk(score, dim=0, $k$)
23             key_recall = concat(
24               past_key_values[$i$].key[$l_{\text{global}}$:],
25               past_key_values[$i$].key[idx_recall],
26               past_key_values[$i$].key[:-$l_{\text{local}}$]
27             )
28             value_recall = concat(
29               past_key_values[$i$].value[$l_{\text{global}}$:],
30               past_key_values[$i$].value[idx_recall],
31               past_key_values[$i$].value[:-$l_{\text{local}}$]
32             )
33             $\triangleright$ Compute self-attention with position embedding
34             attention_output = rope_attention(query, key_recall, value_recall)
35             hidden_states = ffn(attention_output)
36
37      $\triangleright$ Move to the next chunk
38       $s=e$
39       $e=e+l_{\text{chunk}}$
40
return hidden_states, past_key_values
Algorithm 1 Prefilling Phase
Input: KV cache past_key_values, the last token index last_token, number of tokens $L$, the number of starting tokens $l_{\text{global}}$, the number of ending tokens $l_{\text{local}}$.
1 hidden_states = embedding(last_token)
2 $\triangleright$ Forward through each transformer layer
3 for $i\leftarrow 0$ to $N_{\text{layer}}-1$ do
4       $\triangleright$ Project to Q, K, V
5       query = q_proj(hidden_states)
6       key = k_proj(hidden_states)
7       value = v_proj(hidden_states)
8       $\triangleright$ Update KV cache for this layer
9       key = concat(past_key_values[$i$].key, key)
10       value = concat(past_key_values[$i$].value, value)
11       past_key_values[$i$] = (key, value)
12       $\triangleright$ Position-agnostic cache selection
13       score = query @ key[$l_{\text{global}}$:-$l_{\text{local}}$]${}^{\texttt{T}}$
14       idx_recall = topk(score, dim=0, $k$)
15       key_recall = concat(
16         past_key_values[$i$].key[$l_{\text{global}}$:],
17         past_key_values[$i$].key[idx_recall],
18         past_key_values[$i$].key[:-$l_{\text{local}}$]
19       )
20       value_recall = concat(
21         past_key_values[$i$].value[$l_{\text{global}}$:],
22         past_key_values[$i$].value[idx_recall],
23         past_key_values[$i$].value[:-$l_{\text{local}}$]
24       )
25       $\triangleright$ Compute self-attention with position embedding
26       attention_output = rope_attention(query, key_recall, value_recall)
27       hidden_states = ffn(attention_output)
28
29$\triangleright$ Get the next token prediction
30 logits = lm_head(hidden_states)
31 next_token = argmax(logits)
return next_token, past_key_values
Algorithm 2 Decoding Phase
Appendix B More Validation on Effectiveness
@@
S-Doc M-Doc Sum ICL Syn. Code Avg. Close- Ended Open- Ended Avg.
LLaMA3-8B-8K 18.08 13.17 21.42 63.95 10.45 69.89 35.48 46.11 2.79 17.23
+ StreamingLLM 15.14 8.90 18.36 58.19 7.28 70.27 32.49 36.52 1.36 13.08
+ ReAttention 19.08 11.33 20.69 63.40 8.95 70.38 35.03 44.94 2.43 16.60
LLaMA3.1-8B-128K 23.61 15.29 16.15 64.61 31.81 70.63 38.79 45.30 0.50 15.44
+ StreamingLLM 21.67 11.46 14.01 61.05 33.43 70.50 37.06 42.36 0.38 14.37
+ ReAttention 23.97 15.22 15.66 64.54 31.20 70.48 38.63 45.25 0.50 15.42
LLaMA3.2-3B-128K 34.45 26.24 24.46 59.29 40.11 55.73 40.76 55.75 20.63 32.33
+ StreamingLLM 28.05 19.41 22.67 55.23 22.63 54.16 35.08 43.27 18.66 26.87
+ ReAttention 32.53 24.39 24.13 58.13 37.39 55.17 39.43 53.09 20.12 31.11
LLaMA3.2-3B-128K 26.46 15.60 21.88 67.39 66.62 74.89 46.02 60.78 0.15 20.36
+ StreamingLLM 23.72 12.14 20.66 65.35 56.36 75.06 43.20 55.32 0.16 18.54
+ ReAttention 27.23 15.38 21.67 67.58 66.35 75.26 46.15 61.11 0.18 20.49
Mistral-v0.3-7B-32K 19.93 11.22 22.32 62.05 20.97 65.13 35.61 41.97 4.33 16.88
+ StreamingLLM 17.60 9.79 20.68 59.32 19.42 64.85 34.01 36.49 4.45 15.13
+ ReAttention 20.20 10.64 22.41 61.57 20.58 65.32 35.48 41.95 4.09 16.71
InternLM2.5-7B-1M 44.18 41.06 23.17 61.85 63.17 52.62 47.27 62.96 24.99 37.64
+ StreamingLLM 41.26 34.46 22.26 59.58 53.67 52.97 44.03 57.97 23.30 34.85
+ ReAttention 43.85 41.10 23.14 62.15 63.67 52.72 47.35 63.47 24.36 37.40
Qwen2-7B-128K 37.75 14.82 26.16 62.81 29.00 67.65 41.34 53.43 15.41 28.08
+ StreamingLLM 35.46 11.22 25.11 61.46 23.00 67.39 39.16 50.07 14.30 26.22
+ ReAttention 37.64 14.39 26.24 62.90 27.00 67.56 41.01 53.18 15.35 27.96
Qwen2-72B-128K 44.61 54.04 26.33 65.50 64.33 71.15 54.71 68.17 25.44 39.68
+ StreamingLLM 40.11 40.24 25.29 64.48 54.83 70.91 50.08 60.43 22.38 35.06
+ ReAttention 44.96 53.20 26.18 66.00 62.67 71.39 54.53 67.31 25.34 39.33
Qwen2-1B-32K 24.07 12.96 23.45 58.14 10.17 57.98 33.21 41.28 15.59 24.16
+ StreamingLLM 23.87 10.13 22.87 56.62 10.00 58.32 32.40 37.46 15.31 22.69
+ ReAttention 24.55 13.38 23.44 58.27 10.17 58.58 33.51 41.00 16.10 24.40
@@
Table 5 includes a detailed comparison of various LLMs including LLaMA (Meta, 2024a; Dubey et al., 2024; Meta, 2024b), Mistral(mistralai, 2024), InternLM (Cai et al., 2024b; InternLM, 2024), and Qwen (Yang et al., 2024a) series in LongBench (Bai et al., 2023) and L-Eval (An et al., 2023).
Appendix C More Comparison on Efficiency
Regarding the comparison between full attention and ReAttention in efficiency, the throughput and TTFT (Time To First Token) are two different measuring dimensions. TTFT measures performance under identical conditions (same batch size and context length). As shown in Table 7, our TTFT results demonstrate that ReAttention achieves latency on par with full attention equipped with FlashAttention2 under these controlled conditions. Using a single A800 GPU, we test our ReAttention approach and the standard implementation of Hugging Face’s Llama3-8B (Meta, 2024b) modeling with FlashAttention2 (Dao, 2023). We keep all the settings unchanged and set the batch size to 1.
ReAttention (ms) Baseline (ms)
32k 3984.66 3592.32
64k 10696.22 9823.82
96k 20139.42 18942.57
128k 32328.20 31953.82
192k 64832.10 64475.56
256k 108240.09 OOM
Table 6: Comparison on TTFT.
ReAttention Baseline
token/s batch token/s batch
32k 43.380 8 40.595 2
64k 24.383 5 20.399 1
96k 17.005 4 OOM
128k 11.683 2 OOM
192k 7.2813 1 OOM
256k 5.9680 1 OOM
Table 7: Comparison on Throughput.
Throughput, on the other hand, evaluates maximum processing capacity when fully utilizing available GPU memory. As shown in Table 7, ReAttention demonstrates higher efficiency by supporting larger batch sizes while maintaining comparable or better token processing rates. Besides, beyond context length 64K, full attention equipped with FlashAttention2 (Dao, 2023) encounters memory limitations, while ReAttention continues working. Using a single A800 GPU, we test our ReAttention approach and full attention with FlashAttention2 (Dao, 2023) with LLaMA3-8B (Meta, 2024b). We keep all the settings unchanged. The generation length is 1024. This performance pattern indicates not a trade-off where we sacrifice one capability for another. Rather, it represents extending context length while maintaining competitive throughput and enabling larger batch sizes.
Appendix D More Related Work
The evolution of token eviction techniques begins with DCP (Anagnostidis et al., 2024), Scissorhands (Liu et al., 2024b), and H2O (Zhang et al., 2024d). These approaches demonstrate that retaining only a small subset of tokens with high attention scores could maintain near-equivalent performance while significantly reducing memory and computational overhead. Then, StreamingLLM (Xiao et al., 2023) advances this concept by identifying that critical tokens always occur in the beginning and end part of the attention scope. By preserving only corresponding tokens, StreamingLLM enables stable model outputs after infinite input lengths, dramatically reducing computational and memory costs. However, this approach substantially compromised long-context performance (Li et al., 2025; Xiao et al., 2024b). Subsequent researches explore mitigation strategies, such as head-adaptive methods like FastGen (Ge et al., 2023) and DuoAttention (Xiao et al., 2024b). Specifically, SnapKV (Li et al., 2025) introduced a more nuanced approach to pruning previous KV caches by querying the above context with later input segments, substantially enhancing long-context performance for token eviction techniques. Building upon these foundations, researchers such as PyramidKV (Cai et al., 2024a), PyramidInfer (Yang et al., 2024b), Gemfilter (Shi et al., 2024), and LongGen (Ge et al., 2024) take layer-wise optimization, maintaining inference effectiveness while improving memory efficiency further. Besides, works like VATP (Guo et al., 2024) and InfiniPot (Kim et al., 2024) expand token pruning metrics beyond traditional QK dot product scoring. Our ReAttention draws inspiration from these token eviction methods, specifically applying the principle of selecting KV cache to context extrapolation, enabling comprehensive context perception throughout the inference process while computationally addressing each attention calculation with a finite context, thereby achieving algorithmically infinite context length.
Appendix E Analysis of Cache Feature
E.1 The Method for Obtaining t-SNE Results